Can someone show me how to calculate liters at 800 bar?

Fort Collins, CO

Calculation

Result

Moles of H2 for your community

(8.952e-4 moles H2/day)(5 days)

0.004476 moles H2[NH5]

Liters of H2 for your community at STP

(4.476e-3mol H2)/1X(22.4L)/1mol H2=

.106624 L

Liters of H2 for your community at 800 bar

Number of railroad tanker cars to hold your compressed H2
Answer #1

You should know that at STP (1 bar, 273K) the volume of 1 mole gas is equal to 22.4 liters.

Here given liters of H2 at 22.40 C =0.106624 liters, then 0.004476 moles occupy 0.106624*0.004476 liters = 0.000477 liters.

Now apllying Boyle's law;

P1 V1 = P2 V2

P1= 1 bar (STP)

P2= 800 bar (given)

V1 =0.000477 liters

So V2= (P1 * V1)/ P2

V2 = (1 * 0.000477)/800

V2 = 5.96 * 10-7 liters.